# Rethinking cloud abstractions for tenant-provider cooperative optimization of AI workloads

### v1

@@ Abstract.
AI workloads, often hosted in multi-tenant cloud environments, require vast computational resources but suffer inefficiencies due to limited tenant-provider coordination. Tenants lack infrastructure insights, while providers lack workload details to optimize tasks like partitioning, scheduling, and fault tolerance. We propose the HarmonAIze project to redefine cloud abstractions, enabling cooperative optimization for improved performance, efficiency, resiliency, and sustainability. This paper outlines key opportunities, challenges, and a research agenda to realize this vision.
@@ 1. Introduction
The root cause of these inefficiencies lies in the lack of abstractions enabling cooperative optimization between tenants and providers (§ 2), a challenge shared across AI and other cloud applications (Huang et al., 2024; Bilal et al., 2023; Chen et al., 2022). Such abstractions would need to provide provide semantic information necessary for cross-layer control, allowing tenants and providers to jointly optimize performance, efficiency, and resilience, achieving outcomes that neither could accomplish alone.
To address this gap, we propose widening traditional cloud abstractions to enable tenant-provider collaboration. Specifically, these abstractions expand cloud interface boundaries to provide tenants with additional information typically unavailable to them and enable providers to optimize workloads more effectively. Fig. 1 outlines the architecture of our proposal, HarmonAIze, and illustrates interactions between tenants and providers to enable cross-layer optimization opportunities (§ 3) via micro- and macro-level abstractions (§ 4). Compared to the status quo, HarmonAIze introduces two key innovations: (1) tenants’ control loops are informed by infrastructure-level updates, and (2) providers’ control loops are guided by workload-specific requirements, integrating optimizations that tenants could only achieve with perfect infrastructure visibility in an ideal scenario.
Recent work has introduced ad hoc approaches for optimizing AI workloads in specific cases (Wu et al., 2024b; Mai et al., 2020; Duan et al., 2024). In contrast, HarmonAIze advocates a general framework for cooperative optimization. Workload Intelligence (Huang et al., 2024) proposes dynamic, bi-directional communication for a generic cloud cooperation framework. While HarmonAIze shares some goals with WI, it departs from preserving traditional VM abstractions and VM-oriented mechanisms. Instead, it elevates its abstractions closer to the workload and refactors functionality to leverage unique workloads characteristics, such as their accelerator-driven, iterative nature, performance predictability, proximity to the network, and tight synchronization.
@@
2. AI workloads in the cloud
Today, tenants have multiple options for running AI workloads on cloud providers, depending on their specific requirements, levels of expertise, and available resources. Broadly, tenants can choose from managed Machine Learning as a Service (MLaaS) platforms, use Infrastructure as a Service (IaaS) to set up customized environments, or opt for hybrid configurations that combine on-premises and cloud resources.
MLaaS platforms such as Google Vertex AI (Google Cloud, 2024a) and AWS SageMaker (Amazon Web Services, 2024) offer managed services that streamline large-scale ML pipelines. These platforms provide two primary options: (1) AutoML tools, which deliver a code-free experience but limit customization over model architecture and infrastructure, and (2) custom training interfaces that allow tenants to use ML frameworks of their choice (e.g., PyTorch or TensorFlow), run jobs, and utilize select optimizations like Google’s Reduction Server (Google Cloud, 2024b). However, despite these advantages, MLaaS platforms lack the cooperative mechanisms needed for deeper collaboration between tenants and providers, which we envision as essential for effective workload optimization. Additionally, proprietary APIs reinforce vendor lock-in, restricting flexibility and making it costly for tenants to migrate to other platforms.
For greater control and flexibility, tenants may opt for IaaS solutions, provisioning cloud infrastructure to create tailored environments for AI workloads. Using IaaS, tenants can deploy their own ML pipelines, libraries and containers on virtual machines, attach high-performance GPUs or TPUs as needed, and orchestrate resources with Kubernetes clusters. This approach is suitable for users who need specific configurations, such as optimized distributed ML frameworks or customized distributed execution strategies. However, IaaS solutions hinder cooperative optimization as current deployments do not expose sufficient semantic information about AI workloads to enable cross-layer optimization. Thus, tenants must manually handle resource configuration, scheduling, and optimize ML pipelines — tasks that could be optimized more effectively with cross-layer cooperation.
3. Cross-layer opportunities
To make our discussion concrete, we illustrate several cross-layer optimization opportunities that could benefit from the proposed abstractions. These opportunities span various aspects of AI workload management, including communication, scheduling, fault tolerance, and resource scaling.
O5-Reliability: Resilience in large-scale training is a critical challenge (Tay, 2024). Failures cause delays and often require manual intervention, as tenants lack infrastructure visibility. Tenant-led systems, relying on redundancy, can incur unsustainable overheads (He et al., 2024). This highlights the need for efficient checkpointing (Wang et al., 2023, 2024; Cai et al., 2024; Maurya et al., 2024; Gupta et al., 2024), which could benefit from deeper infrastructure visibility. Providers are well-positioned to manage fault tolerance holistically, avoiding the inefficiencies of treating training tasks as isolated black-boxes.
O6-Energy management: Energy consumption is a critical concern for AI workloads (Wu et al., 2024a). Achieving efficiency requires aligning workload characteristics with infrastructure capabilities (Chung et al., 2024; You et al., 2023). For example, dynamic inference cluster reconfiguration can optimize energy and cost while meeting performance SLOs (Stojkovic et al., 2024). However, such optimizations demand insights into workload needs and infrastructure (e.g., assessing carbon intensity), which are often siloed. Cooperative optimization bridges this gap, enabling better strategies.
Recent works (Table 1) show gains in these areas (O1-6), with cross-layer cooperation expected to amplify the benefits. Beyond these examples, other potential areas for tenant-provider cooperation include profiling and troubleshooting tools, adaptive workload management (Patke et al., 2024), and adaptive communication compression (Abdelmoniem and Canini, 2021; Xin et al., 2023; Markov et al., 2024; Chen et al., 2024).
4. HarmonAIze Overview
@@
As discussed in the previous section, the lack of abstractions to enable cooperative optimization between tenants and providers significantly hinders the efficient execution of AI workloads in cloud environments. To address this challenge, we propose HarmonAIze, a set of abstractions designed to foster cooperative optimization between tenants and providers. These abstractions provide the semantic information necessary for intelligent, cross-layer control, while avoiding leaking sensitive infrastructure details to tenants.
@@
Given this high-level structuring of HarmonAIze, the next critical questions to address are: (1) what abstractions are necessary to enable these cooperative control loops, and (2) how can they be realized in practice while minimizing adoption overhead? In the following section, we present preliminary approaches to tackle these challenges.
@@
| O6 | 53% energy, 38% operational carbon emissions (Stojkovic et al., 2024) | Joint device-workload tuning | Train, Inf |
5. Redefined cloud abstractions
@@
Macro-level abstractions: While maximizing API compatibility is critical for micro-level abstractions, macro-level abstractions allow for a more flexible, clean-slate approach to interface design. The primary goal of macro-level abstractions is to provide tenants with an interface that is both expressive enough to convey meaningful information and narrow enough to protect sensitive infrastructure details. One potential design is for providers to expose a pub/sub interface, enabling tenants to subscribe to infrastructure-level events and updates. These events might include changes in resource availability, network congestion notifications, or failure alerts, empowering tenants to make informed, strategic decisions at the macro level. To ensure broad applicability, this interface should be standardized across providers.
@@ 7. Discussion
Adoption. The adoption of HarmonAIze is underpinned by a natural alignment of incentives between tenants and providers. Tenants gain improved performance, efficiency, and resilience for their workloads, while providers benefit from better infrastructure utilization and the ability to offer differentiated, value-added services. Moreover, a buy-in approach affords incremental deployability. Tenants can adopt HarmonAIze’s abstractions at their own pace, without disrupting existing workloads. Providers, in turn, can continue handling workloads from non-adopting tenants using current practices while leveraging the additional insights provided by HarmonAIze tenants. This gradual and flexible deployment model ensures a smooth transition toward cooperative optimization, minimizing adoption barriers for both tenants and providers.
Timing and enablers. The timing of HarmonAIze’s introduction is critical. If introduced too early, it risks stifling innovation by imposing abstractions on a still-evolving ecosystem. Conversely, if introduced too late, the accumulated technical debt – an issue particularly pronounced in ML systems (Sculley et al., 2015) – and the complexity of standardizing entrenched systems may diminish its impact. In recent years, the pace of innovation in AI and workload deployment has been relentless. However, the ecosystem has begun to coalesce around pre-canned recipes and best practices, with systems and libraries like vLLM (Kwon et al., 2023), DeepSpeed (Rajbhandari et al., 2020), Megatron-LM (Shoeybi et al., 2020), NCCL, and Huggingface’s Transformers (Wolf et al., 2020) becoming de-facto standards. This growing stability signals that the ecosystem is now ripe for a new layer of abstraction — one that enables the cooperative optimizations that HarmonAIze envisions.
@@
Early efforts could then focus on the opportunities identified in § 3 and establish a framework for evaluating the impact of HarmonAIze’s abstractions on representative AI workloads. Such a framework would enable systematic assessment of the performance, efficiency, and resilience gains achievable through cooperative optimization. Using published results, Table 1 illustrates select benefits that are representative of cooperative optimizations that we expect from HarmonAIze.
@@ Acknowledgements.
We are grateful to Amin Vahdat for his insightful feedback and suggestions.References
@@
- Cai et al. (2024) Weilin Cai, Le Qin, and Jiayi Huang. 2024. MoC-System: Efficient Fault Tolerance for Sparse Mixture-of-Experts Model Training. arXiv:2408.04307 [cs.DC]
@@
- Stojkovic et al. (2024) Jovan Stojkovic, Chaojie Zhang, Íñigo Goiri, Josep Torrellas, and Esha Choukse. 2024. DynamoLLM: Designing LLM Inference Clusters for Performance and Energy Efficiency. arXiv:2408.00741 [cs.AI]

### v2

# Cloud abstractions for AI workloads

@@ Abstract.
AI workloads, often hosted in multi-tenant cloud environments, require vast computational resources but suffer inefficiencies due to limited tenant-provider coordination. Tenants lack infrastructure insights, while providers lack workload details to optimize tasks like partitioning, scheduling, and fault tolerance. We propose HarmonAIze to redefine cloud abstractions, enabling cooperative optimization for improved performance, efficiency, resiliency, and sustainability. We outline key opportunities and challenges this vision faces.
@@ 1. Introduction
The root cause of these inefficiencies lies in the lack of abstractions enabling cooperative optimization between tenants and providers (§ 2), a challenge shared across AI and other cloud applications (Huang et al., 2024; Bilal et al., 2023; Chen et al., 2022). Such abstractions would need to provide semantic information necessary for cross-layer control, allowing tenants and providers to jointly optimize performance, efficiency, availability and resilience, achieving outcomes that neither could accomplish alone.
To address this gap, we propose widening traditional cloud abstractions to enable tenant-provider collaboration. Specifically, these abstractions expand cloud interface boundaries to provide tenants with additional information typically unavailable to them and enable providers to optimize workloads more effectively. Fig. 1 outlines the architecture of our proposal, HarmonAIze, and illustrates interactions between tenants and providers to enable cross-layer optimization opportunities (§ 4) via micro- and macro-level abstractions (§ 3). Compared to the status quo, HarmonAIze introduces two key innovations: (1) tenants’ control loops are informed by infrastructure-level insights (e.g., a shift in network load), and (2) providers’ control loops are guided by workload-specific requirements (e.g., required communication patterns), integrating optimizations that tenants could only achieve with perfect infrastructure visibility in an ideal scenario.
Recent work has introduced ad hoc approaches for optimizing AI workloads in specific cases (Wu et al., 2024b; Mai et al., 2020; Duan et al., 2024). In contrast, HarmonAIze advocates a general framework for cooperative optimization. Workload Intelligence (WI) (Huang et al., 2024) proposes dynamic, bi-directional communication for generic cloud cooperation. While HarmonAIze shares some goals with WI, it departs from preserving traditional VM abstractions and VM-oriented mechanisms. Instead, it elevates its abstractions closer to the workload and refactors functionality to leverage unique workloads characteristics, such as their accelerator-driven, iterative nature, performance predictability, proximity to the network, and tight synchronization.
@@
2. Today’s Cloud Fails AI Workloads
Tenants today can choose from a range of cloud-based options to run AI workloads, including managed Machine Learning as a Service (MLaaS) platforms, Infrastructure as a Service (IaaS), or hybrid configurations. Each model presents trade-offs between abstraction and control, and current systems fall short in enabling effective cross-layer optimization.
MLaaS platforms (e.g., Google Vertex AI (Google Cloud, 2024a), AWS SageMaker (Amazon Web Services, 2024)) offer convenient interfaces for launching and scaling training jobs. However, they provide little to no insight into low-level resource behavior or cluster conditions. For example, systems like Google’s Reduction Server (Google Cloud, 2024b) optimize communication, but tenants cannot influence or adapt those strategies at runtime. Furthermore, AutoML features limit control over model architecture and hyperparameters, while proprietary APIs restrict portability across clouds.
IaaS-based deployments, in contrast, grant tenants full control over infrastructure via virtual machines, Kubernetes clusters, and direct access to GPUs or TPUs. However, this control comes at the cost of responsibility. Tenants must manually manage aspects such as job placement, communication library selection, and failure recovery—tasks that are both complex and decoupled from provider-level information.
These limitations lead to three pervasive problems:
Jobs are static when they should be adaptive: Most training jobs are configured statically, with fixed resource allocations and learning settings. This leads to inefficiencies when network conditions, cluster load, or hardware performance fluctuates mid-run. For instance, systems like KungFu (Mai et al., 2020) and Parcae (Duan et al., 2024) show that adapting batch sizes or optimizer settings dynamically can improve convergence and throughput. Today, such adaptations require extensive custom infrastructure.
Tenants and providers operate in the dark: Tenants choose resources blindly, without visibility into current cluster utilization or topology. Communication libraries like NCCL (NVIDIA, 2024) select algorithms (e.g., ring or tree) at initialization, based on assumptions that may no longer hold as the job progresses. Providers, on the other hand, lack hooks to influence job behavior in response to failures or congestion. This gap prevents runtime adaptation or efficient network-aware execution (Wu et al., 2024b; Cao et al., 2024; Liu et al., 2024a).
Resource allocation is clumsy and manual: Cluster schedulers assign resources based on static tenant requests, even though ideal allocations vary over time. Advanced scheduling systems like Pollux (Qiao et al., 2021) and Gandiva (Chaudhary et al., 2020) demonstrate the benefits of adaptive scaling, but require full control over both workload and infrastructure. Public clouds cannot offer this level of control without proper abstractions.
3. Coordinating AI Workloads with HarmonAIze
@@
The lack of abstractions to enable cooperative optimization between tenants and providers significantly hinders the efficient execution of AI workloads in cloud environments. To address this challenge, we propose HarmonAIze, a set of abstractions designed to foster cooperative optimization between tenants and providers.
@@
Given this high-level structuring of HarmonAIze, the next critical questions to address are: (1) what abstractions are necessary to enable these cooperative control loops, and (2) how can they be realized in practice while minimizing adoption overhead? We now present preliminary approaches to tackle these challenges.
@@
| O6 | 22.61 up time between failures (Xiong et al., 2024) | Joint collection of observability data | Train |
| O7 | 53% energy, 38% operational carbon emissions (Stojkovic et al., 2025) | Joint device-workload tuning | Train, Inf |
3.1. Redefined cloud abstractions
@@
Macro-level abstractions: While maximizing API compatibility is critical for micro-level abstractions, macro-level abstractions allow for a more flexible, clean-slate approach to interface design. The primary goal of macro-level abstractions is to provide tenants with an interface that is both expressive enough to convey meaningful information and narrow enough to protect sensitive infrastructure details. One potential design is for providers to expose a pub/sub interface, enabling tenants to subscribe to infrastructure-level events and updates. For example, if a provider frees up a GPU, it could send a notification to all subscribed tenants. To prevent contention when multiple tenants attempt to allocate the same resource, the system could implement mechanisms such as priority-based allocation, reservation windows, or price-based auctioning (Chaudhary et al., 2020). To ensure broad applicability, this interface should be standardized across providers.
@@ 3.2. Control loops
4. Cross-Layer Benefits of HarmonAIze
We now illustrate how HarmonAIze enables a range of cross-layer optimizations that are difficult or impossible to realize in today’s cloud environments. In each case, HarmonAIze acts as the missing link between tenant-level logic and provider-level infrastructure, enabling both sides to make more informed, coordinated decisions.
O5-Reliability: Resilience in large-scale training is a critical challenge (Tay, 2024). Failures cause delays and often require manual intervention, as tenants lack infrastructure visibility. Tenant-led systems, relying on redundancy, can incur unsustainable overheads (He et al., 2024). This highlights the need for efficient checkpointing (Wang et al., 2023, 2024; Cai et al., 2025; Maurya et al., 2024; Gupta et al., 2024), which could benefit from deeper infrastructure visibility. Providers are well-positioned to manage fault tolerance holistically, avoiding the inefficiencies of treating training tasks as isolated black-boxes.
O6-Availability: Recovery techniques, e.g., check pointing, mitigate the impact of failures, but such failures still cause a significant loss in revenue, e.g., as much as a $30K (Qian et al., 2024) . Recent work (Xiong et al., 2024) shows that the mean time between failures can be increased by 22.61×% if defective components are detected proactively and removed quickly. However, such approaches require knowledge of a tenant’s workloads and events coupled with the provider’s ability to perform proactive measurements using information about their infrastructure. Tenant and provider cooperation ensures that diagnostics framework have access to the application level events and signals needed to disambiguate problems and the infrastructure level observability data needed to triage and identify the root cause of these problems.
O7-Energy management: Energy consumption is a critical concern for AI workloads (Wu et al., 2024a). Achieving efficiency requires aligning workload characteristics with infrastructure capabilities (Chung et al., 2024; You et al., 2023). For example, dynamic inference cluster reconfiguration can optimize energy and cost while meeting performance SLOs (Stojkovic et al., 2025). However, such optimizations demand insights into workload needs and infrastructure (e.g., assessing carbon intensity), which are often siloed. Cooperative optimization bridges this gap, enabling better strategies.
Recent works (Table 1) show gains in these areas (O1-7), with cross-layer cooperation expected to amplify the benefits. Beyond these examples, other potential areas for tenant-provider cooperation include profiling and troubleshooting tools, adaptive workload management (Patke et al., 2024), and adaptive communication compression (Abdelmoniem and Canini, 2021; Xin et al., 2023; Markov et al., 2024; Chen et al., 2024).
@@ 5. Discussion
Adoption. The adoption of HarmonAIze is underpinned by a natural alignment of incentives between tenants and providers. Tenants gain improved performance, efficiency, availability and resilience for their workloads, while providers benefit from better infrastructure utilization and the ability to offer differentiated, value-added services. Moreover, a buy-in approach affords incremental deployability. Tenants can adopt HarmonAIze’s abstractions at their own pace, without disrupting existing workloads. Providers, in turn, can continue handling workloads from non-adopting tenants using current practices while leveraging the additional insights provided by HarmonAIze tenants. This gradual deployment model ensures a smooth transition toward cooperative optimization.
Timing and enablers. The timing of HarmonAIze’s introduction is critical. If introduced too early, it risks stifling innovation by imposing abstractions on a still-evolving ecosystem. Conversely, if introduced too late, the accumulated technical debt – an issue particularly pronounced in ML systems (Sculley et al., 2015) – and the complexity of standardizing entrenched systems may diminish its impact. In recent years, the pace of innovation in AI and workload deployment has been relentless. However, the ecosystem has begun to coalesce around pre-canned recipes and best practices, with systems and libraries like vLLM (Kwon et al., 2023), DeepSpeed (Rajbhandari et al., 2020), Megatron-LM (Shoeybi et al., 2020), NCCL, and Huggingface’s Transformers (Wolf et al., 2020) becoming de-facto standards. This growing stability signals that the ecosystem is now ripe for a new layer of abstraction, as we envision.
@@
Early efforts could then focus on the opportunities identified in § 4 and establish a framework for evaluating the impact of HarmonAIze’s abstractions on representative AI workloads. Such a framework would enable systematic assessment of the performance, efficiency, availability and resilience gains achievable through cooperative optimization. Using published results, Table 1 illustrates select benefits that are representative of cooperative optimizations that we expect from HarmonAIze.
@@ Acknowledgements.
We are grateful to Amin Vahdat for his insightful feedback and suggestions. This work has been partially supported by Knut and Alice Wallenberg Foundation (Wallenberg Scholar Grant for Prof. Dejan Kostić), as well as NSF grant 2212193.References
@@
- Cai et al. (2025) Weilin Cai, Le Qin, and Jiayi Huang. 2025. MoC-System: Efficient Fault Tolerance for Sparse Mixture-of-Experts Model Training. In ASPLOS.
@@
- Qian et al. (2024) Kun Qian, Yongqing Xi, Jiamin Cao, Jiaqi Gao, Yichi Xu, Yu Guan, Binzhang Fu, Xuemei Shi, Fangbo Zhu, Rui Miao, Chao Wang, Peng Wang, Pengcheng Zhang, Xianlong Zeng, Eddie Ruan, Zhiping Yao, Ennan Zhai, and Dennis Cai. 2024. Alibaba HPN: A Data Center Network for Large Language Model Training. In SIGCOMM.
@@
- Stojkovic et al. (2025) Jovan Stojkovic, Chaojie Zhang, Íñigo Goiri, Josep Torrellas, and Esha Choukse. 2025. DynamoLLM: Designing LLM Inference Clusters for Performance and Energy Efficiency. In HPCA.
@@
- Xiong et al. (2024) Yifan Xiong, Yuting Jiang, Ziyue Yang, Lei Qu, Guoshuai Zhao, Shuguang Liu, Dong Zhong, Boris Pinzur, Jie Zhang, Yang Wang, Jithin Jose, Hossein Pourreza, Jeff Baxter, Kushal Datta, Prabhat Ram, Luke Melton, Joe Chau, Peng Cheng, Yongqiang Xiong, and Lidong Zhou. 2024. SuperBench: Improving Cloud AI Infrastructure Reliability with Proactive Validation. In USENIX ATC.